OpenStudy (anonymous):

Find the inverse of the function f(x) = (2x+5)/3

OpenStudy (anonymous):

$$x=\frac{2f^{-1}(x)+5}{3}$$ Now I am sure you can do the rest

OpenStudy (anonymous):

wait, so I'm not sure what to do, I don't understand functions. Do I simplify?

OpenStudy (anonymous):

$$f^{-1}(x)$$ is the inverse function of f, this means that $$f(f^{-1}(x))=x$$

OpenStudy (anonymous):

Substituting $$f^{-1}(x)$$ in for the x appearing in that expression we get

OpenStudy (anonymous):

$$x=\frac{2f^{-1}(x)+5}{3}$$

OpenStudy (anonymous):

so is it x^2 +5 / 3

OpenStudy (anonymous):

Now solving for $$f^{-1}(x)$$ gives us, $$\frac{3x-5}{2}=f^{-1}(x)$$.

OpenStudy (anonymous):

what?? how did you come up with that? this is so confusing

OpenStudy (anonymous):

Try using khan academy or somthing

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

keywords are "functions/Inverse functions"